26 November 2007

RHIC Startup:

Tune injectors up to 1E11 for deuterons and 1E9 for Gold.  Intensity had been dropping throughout the weekend.  It appears to be improved now.

Tune jump in Blue at the start of ramp whenever the replay file is used.   There are three strategies for dealing with this, but all involve removing rapid tune changes at the beginning of the ramp, either by regenerating the ramp file without tune corrections, filtering out the early corrections, or modifying the software to apply the corrections slowly. 

Rebucketing in Blue is required, giving to big longitudinal emittance.  MikeB would like to trouble-shoot the RF to see if that  is the source of the rebucketing problem.  The rebucketing display also fails to show both beams on occasion. 

There appears to be coupling near the end of ramp in Blue.  Yun will look into it today.

Radial excursion around transition in yellow has gotten large again.  MikeB noticed that Blue Orbit appears to show a bump when the Yellow is going through transition.  Blue goes through transition 8 seconds before Yellow, so Angelika thinks the display is correct, and the excursion in Blue is in time with passing through Blue transition and not related to Yellow transition. 

It was agreed that today beginning with the evening shift, the experiments will have collisions from 4:00PM through to 8:00AM, with the day shift going to beam development. 

First try the new replay file, then fix rebucketing.

Last night we attempted more bunches, but had to retreat because of losses at t260. 

Collimator controllers failed over the weekend, causing a beam loss when a controller walked a collimator into the beam over a one and a half minute period.  But the current beam excursion of 1.5mm is too large for collimator tracking.   

PHENIX has a safety issue with their front door.  Fred Kobichuk has called in the work order, but nothing has been done yet.  The entire door frame will probably have to be replaced.  PHENIX wants more beam intensity.  Total beam is down ~2, and vertex efficiency is down a factor 4-5.  Both need fixing.

During a fire alarm in Bldg 1010B (apparently caused by a faulty smoke alarm) Steve Jao made an access to TTB to replace a chip in the Current Transformer amplifier in Section 29.  Earlier attempts to fix this beam current monitor did not respond to any of the easy fixes.

In response to a request from the vacuum group, scrubbing will not be done until it is required.  And currently it is not required, so there is no plan for scrubbing.

Alex wants us to reduce the gap volts in the Band 2 system.  This applies to both deuterons and gold.  

Maintenance day planning: PHENIX has a light day's work on power supplies.  PHENIX also has a non-functioning pulser in the South Tunnel that needs repair by Steve Boose.  Also some routine Air Conditioning maintenance that Charlie Pearson is doing.  Charlies' folks are requested to check in with PHENIX Control Room before making access and when done.  STAR says they will use the entire 8 hour access for unspecified work.  There is an access request for the 12OClock Hall to replace a pump on the JET.

We have 56x56 with bunch intensities of 1E11 Blue and 0.6E9 in Yellow with good rebucketing (the best so far).  Landau cavities have been retimed. 

Tonight we will fit in three 7 hour stores between 3:00PM and 9:30AM tomorrow morning.

APEX begins on Wednesday 5 Dec 2007.  Since the plan is to do a maintenance day the same day, APEX will start at 00:01 AM.


Scheduling Physicist: Mike Sivertz Last updated:
Email:sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020